Terry’s utility function over leisure (L) and other goods (Y ) is U(L, Y ) = Y + LY. The associated marginal utilities are MUY = 1 + L and MUL = Y. He purchases other goods at a price of $1, out of the income he earns from working. Show that, no matter what Terry’s wage rate, the optimal number of hours of leisure that he consumes is always the same.

(a)  What is the number of hours he would like to have for leisure?

(b)  Determine the MRS of leisure for labour

(c)  Draw a leisure-influenced labor curve


1
Expert's answer
2021-05-05T13:34:09-0400

a)

"U(L,Y)=Y+LY"

"MU_Y=1+L"

"MU_L=Y"

"P_y=1"

Budget line: "WL+Y=24W.......................(1)"

At optimality we know

"MRS=\frac{MU_Y}{MU_L}=\frac{1+L}{Y}=\frac{P_y}{P_l}=\frac{1}{w}"


"\frac{Y}{1+L}=w"


"Y=w(1+L)......................" put in (1)

"WL+W(1+L)=24W"

"WL+W+WL=24W"

"2WL=23W"

"L=\frac {23W}{2W}=11.5"


leisure=11.5

"\Delta Y=W(12.5)"


b)

"MRS=\frac{1_l}{Y}=\frac{12.5}{Y}=\frac{12.5}{12.5(W)}=\frac{1}{W}"

"MRS=\frac{1}{W}"


c)

leisure remain same irrespective of wage rate

Thus labor also remains same